OpenStudy (charlotte123):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

jimthompson5910 (jim_thompson5910):

hint: your answer will have a slope of -1 since y = -x-4 has a slope of -1 parallel lines have the same slope

OpenStudy (charlotte123):

@jim_thompson5910 so would it be B? :O

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

y = -x - 2 1 = -(-3) - 2... plug in (-3,1) 1 = 3 - 2 1 = 1 So this shows that (-3, 1) is on the line y = -x-2

OpenStudy (charlotte123):

thanks Jim :D

jimthompson5910 (jim_thompson5910):

np